WORK
3)Two executives in cities 400 miles apart drive to a business meeting at a location on the line between their cities. They meet after 4 hrs. Find the speed of each car if one car travels 20 miles per hour faster than the other does.



Thank you in advance! Hoping for your immediate reply.

Answer by sudhanshu_kmr(1152)


assume speed of one car = x miles per hour
then speed of another = x+20 miles per hour
then 4(x + x+20) = 400
=> 8x+ 80 = 400
=> 8x = 320

=> x = 320/8= 40
so, speed of cars ... 40 and 60